Pilar spent 3/8 of her salary and had Php 5,250.00 left. What is her monthly salary?

Answer by stanbon(75887)

Pilar spent 3/8 of her salary and had Php 5,250.00 left.
What is her monthly salary?
-----
Let her monthly salary be "x"::
x - (3/8)x = 5250
-----
(5/8)x = 5250
----
x = (8/5)5250
-----
x = 8*1050
x = Php 8,400
--------------
Cheers,
Stan H.
-------------